OpenStudy (anonymous):

Evaluate the intregral: cosx/9+sin^2x dx

OpenStudy (mimi_x3):

\[\large \int\limits\frac{cosx}{9+\sin^{2}x} dx\] It's a u-subsitution. let u= sinx => du/cosx

OpenStudy (anonymous):

thanks!